You are conducting a study to see if the proportion of women over 40 who regularly have mammograms is significantly different from 79%. With H1 : p ≠ 79% you obtain a test statistic of z=2.321. Find the p-value accurate to 4 decimal places.
Answer #1

This is a two tailed test. Hence,

p - Value = 2*P(z > 2.321) = 2*0.01014 = 0.0203